6. Freds Junk Shop is expecting to pay unusual dividends for the next two years, then Fred level off at 3%. Freds next two
Answer #1

Dividend in Year 1, D1 = $1.89
Dividend in Year 2, D2 = $2.54

Growth Rate, g = 3%
Required Return, rs = 6%

Dividend in Year 3, D3 = D2 * (1 + g)
Dividend in Year 3, D3 = $2.54 * 1.03
Dividend in Year 3, D3 = $2.6162

Price in two years, P2 = D3 / (rs - g)
Price in two years, P2 = $2.6162 / (0.06 - 0.03)
Price in two years, P2 = $2.6162 / 0.03
Price in two years, P2 = $87.21

Price today, P0 = D1 / (1 + rs) + D2 / (1 + rs)^2 + P2 / (1 + rs)^2
Price today, P0 = $1.89 / 1.06 + $2.54 / 1.06^2 + $87.21 / 1.06^2
Price today, P0 = $81.66

Dividend in Year 12, D12 = D2 * (1 + g)^10
Dividend in Year 12, D12 = $2.54 * 1.03^10
Dividend in Year 12, D12 = $3.41355

Price in eleven years, P11 = D12 / (rs - g)
Price in eleven years, P11 = $3.41355 / (0.06 - 0.03)
Price in eleven years, P11 = $3.41355 / 0.03
Price in eleven years, P11 = $113.78

Value of a share of FJS today (P0) is $81.66
Value of a share of FJS in two years (P2) is $87.21
Value of a share of FJS in eleven years (P11) is $113.78